Can you help me solve this 3x3 equation:
2x+y-z=3
2x+2y-z=-1
-x-y+z=4

Found 2 solutions by richwmiller, stanbon:
Answer by richwmiller(17219)

no similar problems.
You have already submitted others like this today.

Answer by stanbon(75887)

2x +y -z =3
2x +2y -z =-1
-x -y +z =4
---------------
Make the 3rd row the 1st row:
-----------------
-x -y -z = 3
2x +y -z = 3
2x +2y -z = -1
--------
Add 2 times the 1st row to the 2nd and to the 3rd rows.
-x-y +z = 3
0 -y -3z = 9
0 +0 -3z = 5
-------
Solve the 3rd row for "z":
z = -5/3
--------
Substitute into the 2nd row and solve for "y":
-y -3*(-5/3) = 9
-y + 5 = 9
y = -4
--------------
Substitute into the 1st row and solve for "x":
-x -(-4) -(-5/3) = 3
-x + 12/3 +(5/3) = 3
-x + (17/3) = (9/3)
x = 8/3
=================
Cheers,
Stan H.
================
============================
--------